# Purification of Noisy Measurements and Faithful Distillation of Entanglement

### v1

@@ Abstract
We first apply the purification of a noisy measurement to enable entanglement distillation and then show the robustness of the protocol by considering noise that may be contained in CNOT gates. The purification protocol will attain the entanglement distillability from two-qubit states. For a pair of isotropic states obtained after twirling shared states, two parties perform bilateral CNOT gates followed by noisy measurements. Alice and Bob exploit $n-1$ and $m-1$ additional qubits, respectively, for purifying noisy measurements. A resulting singlet fidelity of Alice and Bob relies on measurements on $n$ ad $m$ qubits and is given by,
$F^{\prime}=\frac{F^{2}+\left(\frac{1-F}{3}\right)^{2}+g^{(n,m)}(F)}{F^{2}+2F\left(\frac{1-F}{3}\right)+5\left(\frac{1-F}{3}\right)^{2}+4g^{(n,m)}(F)}$
@@
As $n$ and $m$ tends to be large, we have that $g^{(n,m)}(F)$ converges to zero and reproduce the fidelity in Eq. (12) of a noiseless instance.
Figure 3: In a noiseless case, two-qubit states having a singlet fraction $F>1/2$ are distillable. With noisy measurements with a noise fraction $p=0.1$, the distillation protocol works for entangled states with a singlet fraction $F>0.617$. The purification of noisy measurements leads to an improvement: entanglement can be distilled from entangled states with $F>0.505$. The protocol is also robust: even for measurements having different error rates $p_{i}\in(0.025,0.175)$, see Eq. (16), the protocol faithfully increases a singlet fidelity (grey).
For the full generality in error rates on the measurements, we suppose that error rates in independent individual measurements are not identical. Let $p_{k}^{A}$ denote a noise fraction of the $k$-the measurement of Alice and $p_{k}^{B}$ for Bob, see Eq. (2). The resulting fidelity is obtained with $g^{(n,m)}(F)$ in Eq. (13) defined by functions $r_{even}^{(n,m)}$ and $r_{odd}^{(n,m)}$,
@@
We detail the derivation in Appendix A. In Fig. 3, we consider an instance with $p_{i}\in(0.025,0.175)$ and show that entanglement increases faithfully.
Then, the distillation protocol increases entanglement whenever $F^{\prime}>F$, for which we compute:
@@
Thus, we have $F^{\prime}>F$ for $F\in(L^{(n,m)},1)$ where the lower bound $L^{(n,m)}$ depends on the numbers of measurement outcomes $n$ and $m$ for purifying measurements and a noise fraction $p$ but not an initial fidelity $F$. Note that $L^{(n,m)}$ converges to $1/2$ as $n$ and $m$ tend to be large; hence, the noiseless instance is reproduced.
Examples. We consider entanglement distillation with noisy measurements in a realistic scenario. We are interested in the lower bound $L^{(n,m)}$ to see how the purification can push the limitations in distillability.
For noisy measurements with $p=0.2$, i.e., an error rate $10\%$, we have lower bounds as follows, n 1 2 3 4 $L^{(n,1)}$ 0.781 0.640 0.627 0.625 $L^{(n,2)}$ 0.640 0.525 0.514 0.513 $L^{(n,3)}$ 0.627 0.514 0.503 0.502 From $L^{(n,1)}$, we conclude the purification on noisy measurements of a single party yet achieves a limited range: it improves up to $0.625$ by the purification of noisy measurements of Alice. We also find that from $L^{(n,2)}$ and $L^{(n,3)}$, two additional qubits, i.e., $(n,m)=(3,3)$, are cost-effective as the lower bound is immediately reach $0.503$.
Then, we consider measurements with a noise fraction $p=0.1$, meaning that an error rate is $5\%$, we have lower bounds for distillability as follows, n 1 2 3 4 $L^{(n,1)}$ 0.617 0.559 0.556 0.556 $L^{(n,2)}$ 0.559 0.505 0.500 0.500 which shows that the purification of noisy measurements with single additional qubits on two parties nearly recovers the noiseless bound $1/2$, see Fig. 3 in which the improvement of lower bounds is shown.
Robustness against noise. For the robustness of the purification protocol, we consider noise in a collective CNOT gate $V_{n}$, which may be decomposed as
@@
Since a quantum operation is implemented by single-qubit and CNOT operations in practice, we consider a noise fraction $\epsilon$ in a CNOT gate [41],
@@
Then, a measurement is described by $r_{i}^{(n)}$ for $i=0,1$ in Eq. (8) that can be expressed by a recurrence relation,

### v3

@@ III.2 Noisy scenario
It thus suffices to consider noisy measurements with depolarization noise to show that the purification of noisy measurements improves entanglement distillation. For a pair of isotropic states obtained after twirling shared states, two parties perform bilateral CNOT gates followed by noisy measurements. Alice and Bob exploit $n-1$ and $m-1$ additional qubits, respectively, for purifying noisy measurements. A resulting singlet fidelity of Alice and Bob relies on measurements on $n$ and $m$ qubits and is computed and expressed as,
$F^{\prime}=\frac{F^{2}+\left(\frac{1-F}{3}\right)^{2}+g^{(n,m)}(F)}{F^{2}+2F\left(\frac{1-F}{3}\right)+5\left(\frac{1-F}{3}\right)^{2}+4g^{(n,m)}(F)}$ (21)
@@
For large $n$ and $m$, we have $g^{(n,m)}(F)$ converging to zero and reproduce the fidelity in Eq. (19) of a noiseless instance. Note that the probability of having identical outcomes in the distillation protocol is given by
$p_{succ}^{(n,m)}(F)$ $=$ $r_{even}^{(n,m)}\left(F^{2}+2F\left(\frac{1-F}{3}\right)+5\left(\frac{1-F}{3}\right)^{2}\right)$ (24)
$+r_{odd}^{(n,m)}\left(4F\left(\frac{1-F}{3}\right)+4\left(\frac{1-F}{3}\right)^{2}\right).$
In Supplemental Material, we detail in relation to the probability above and the fidelities in Eq. (21), see Table 2, where one additional qubit is cost-effective.
Figure 4: A measurement error rate, given by $s$, see also Eq. (14), does not single out a POVM element, which can be characterized by a parameter $t\in[-\alpha,\alpha]$ where $\alpha=\sqrt{s(1-s)}$. For a measurement error $5\%$, i.e., $s=0.05$, we have $\alpha\approx 0.2179$. The worst case that maximizes the lower bound for distillability is when $t=0$, see Appendix C where it is shown that the lower bound decreases as $|t|$ increases. (a) A depolarization noise on measurements, $t=0$, leads to the worst-case scenario in that a lower bound for distillability is the highest. The lower bound for distillability is found $L=0.617$ (red). The purification protocol with a single target qubit ($n=m=2$) improves the lower bound up to $L=0.505$ (yellow). (b) Noisy measurements with $t=\alpha$ in Eq. (14) show a lower bound for distillability $L=0.5$ (purple). Hence, a noisy measurement can be used to distill entangled states having a fidelity $F>1/2$. In this case, the purification of noisy measurements can improve the rate of distilling entanglement (blue).
For the full generality, we suppose that error rates in independent individual measurements are not identical. Let $p_{k}^{A}$ denote a noise fraction of the $k$-th measurement of Alice and $p_{k}^{B}$ for Bob, see Eq. (3). The resulting fidelity is obtained with $g^{(n,m)}(F)$ in Eq. (22) defined by functions $r_{even}^{(n,m)}$ and $r_{odd}^{(n,m)}$,
@@
We detail the derivation in Supplemental Material. In Fig. 3, we consider an instance with $p_{i}\in(0.025,0.175)$ and show that entanglement increases faithfully.
Then, the distillation protocol increases entanglement whenever $F^{\prime}>F$:
@@
We have $F^{\prime}>F$ for $F\in(L^{(n,m)},1)$ where the lower bound $L^{(n,m)}$ for the distillability depends on the numbers of measurements $n$ and $m$ for purifying measurements and a noise fraction $p$. Note that $L^{(n,m)}$ converges to $1/2$ as $n$ and $m$ tend to be large; hence, the noiseless instance is reproduced.
In a realistic scenario with an error rate $5\%$, we have a cost-effective and resource-efficient strategy $(n,m)=(2,2)$, i.e., single target qubits. A lower bound $L^{(n,m)}$ is computed in Table 1.
n 1 2 3 4
$L^{(n,1)}$ 0.617 0.559 0.556 0.556
$L^{(n,2)}$ 0.559 0.505 0.500 0.500
Table 1: The lower bound of fidelity at which entanglement can be distilled. $L^{(n,m)}$ indicates that Alice performs $n$ measurements, while Bob performs $m$ measurements.
One can find that the purification of noisy measurements with one additional qubit on each party recovers the noiseless bound $1/2$, see Fig. 3.
III.3 Robustness
We detail the analysis when CNOT gates in the purification of noisy measurements are noisy. We consider noise in a collective CNOT gate $V_{n}$, which may be decomposed as
@@ III.3 Robustness
Since a quantum operation is implemented by single-qubit and CNOT operations in practice, we consider a noise fraction $\epsilon$ in a CNOT gate
@@
We here derive a lower bound $L_{\epsilon}^{(n,m)}$ such that $F^{{}^{\prime}}>F$ for $F>L_{\epsilon}^{(n,m)}$, where the lower bound relies on an error rate $\epsilon$. In particular, we show that $L_{\epsilon}^{(n,m)}$ reaches an $\epsilon$-close bound to $1/2$. It suffices to consider that $n=m$. A lower bound found is given by,
$r_{0}^{(n+1)}$ $=$ $(1-\epsilon)r_{0}^{(n)}(1-\frac{p}{2})+\frac{\epsilon}{4}(r_{0}^{(n)}+r_{1}^{(n)})~{}~{}\mathrm{and}~{}~{}$
@@
From the relation above, we write by $u$ the limit point,
$u=\lim_{n\rightarrow\infty}\frac{r_{1}^{(n)}}{r_{0}^{(n)}}.$ (30)
It follows that,
$u=\frac{(1-\epsilon)u(\frac{p}{2})+\frac{\epsilon}{4}(1+u)}{(1-\epsilon)(1-\frac{p}{2})+\frac{\epsilon}{4}(1+u)},~{}~{}\mathrm{from~{}which~{}}~{}u=2(1-p)(1-\frac{1}{\epsilon})+\sqrt{5-4p(2-p)+\frac{4(1-p)^{2}}{\epsilon}(\frac{1}{\epsilon}-2)}.$
Consequently, when $n$ tends to be large, the lower bound converges $L_{\epsilon}^{(\infty)}$ as follows,
$L_{\epsilon}^{(\infty)}=\lim_{n\rightarrow\infty}L_{\epsilon}^{(n)}=\frac{2(1-p)^{2}(1-2\epsilon)+\epsilon^{2}(3-2(2-p)p)+\epsilon\sqrt{4(1-p)^{2}(1-2\epsilon)+\epsilon^{2}(5+4(-2+p)p)}}{4(1-\epsilon)^{2}(1-p)^{2}}.$ (31)
Note that the lower bound above approaches continuously to $1/2$ as $\epsilon$ converges to 00.
Figure 5: When a CNOT gate contains an error fraction $\epsilon$, see Eq. (29), the purification of noisy measurements recovers the lower bound for distilling entanglement up to $L_{\epsilon}^{(\infty)}$ in Eq. (31), which may be strictly larger than $1/2$. The lower bound $L_{\epsilon}^{(\infty)}$ increases as a noise fraction $\epsilon$ increases.